OpenStudy (stephanos100):

PLS HELPP!!!! TOKENS! Find all polar coordinates of point P where P = (7, pi divided by 3)

OpenStudy (anonymous):

\[P=\left( 7,\frac{ \pi }{ 3 }+2n \pi \right)\],where n is an integer.

OpenStudy (anonymous):

\[P=\left( -7,\frac{ \pi }{ 3 }\pm \pi \right)\]

jimthompson5910 (jim_thompson5910):

@surjithayer your second expression is a bit off

jimthompson5910 (jim_thompson5910):

It should be \[\Large P=\left( -7,\frac{ \pi }{ 3 }+2n \pi +\pi\right)\] \[\Large P=\left( -7,\frac{ 4\pi }{ 3 }+2n \pi\right)\]

OpenStudy (anonymous):

yes you are correct.

OpenStudy (stephanos100):

these are my options though: (7, pi divided by 3 + 2nπ) or (-7, pi divided by 3 + 2nπ) (7, pi divided by 3 + 2nπ) or (-7, pi divided by 3 + (2n + 1)π) (7, pi divided by 3 + (2n + 1)π) or (-7, pi divided by 3 + 2nπ) (7, pi divided by 3 + nπ) or (-7, pi divided by 3 + nπ)

OpenStudy (stephanos100):

@jim_thompson5910

OpenStudy (anonymous):

(7, pi divided by 3 + 2nπ) or (-7, pi divided by 3 + (2n + 1)π)